Suppose x is a normally distributed random variable with muequals12 and sigmaequals2. Find each of the following probabilities. a.​ P(xgreater than or equals15​)   b.​ P(xless than or equals8​)   c.​ P(12.52less than or equalsxless than or equals16.74​)   d.​ P(7.6less than or equalsxless than or equals15.1​)
Answer #1

a)

for normal distribution z score =(X-μ)/σ
here mean=       μ= 12
std deviation   =σ= 2.0000

P(xgreater than or equals15​) :

probability = P(X>15) = P(Z>1.5)= 1-P(Z<1.5)= 1-0.9332= 0.0668

b)

P(xless than or equals8​) :

probability = P(X<8) = P(Z<-2)= 0.0228

c)

P(12.52less than or equalsxless than or equals16.74​) :

probability = P(12.52<X<16.74) = P(0.26<Z<2.37)= 0.9911-0.6026= 0.3885

d)

probability = P(7.6<X<15.1) = P(-2.2<Z<1.55)= 0.9394-0.0139= 0.9255